OpenStudy (anonymous):

How many atoms of calcium in 1.76 moles

OpenStudy (accessdenied):

The term 'mole' is a lot like the term 'dozen,' like talking about dozens of apples. There are 12 apples in a dozen. In 3 dozens? 3*12 = 36 apples! We just multiply the quantity of dozens by the number of apples per dozen (12). There are 6.022x10^23 atoms in a mole. In 3 moles? 3x6.022x10^23 = 18.07x10^23 = 1.807x10^24 atoms! We just multiply the quantity of moles by the number of atoms per mole (avogadro's number, or 6.022x10^23). Can you see how this is done with your problem? (In 1.76 moles)